### v1

# Federated Continual Graph Learning

Abstract
In the era of big data, managing evolving graph data poses substantial challenges due to storage costs and privacy issues. Training graph neural networks (GNNs) on such evolving data usually causes catastrophic forgetting, impairing performance on earlier tasks. Despite existing continual graph learning (CGL) methods mitigating this to some extent, they predominantly operate in centralized architectures and overlook the potential of distributed graph databases to harness collective intelligence for enhanced performance optimization. To address these challenges, we present a pioneering study on Federated Continual Graph Learning (FCGL), which adapts GNNs to multiple evolving graphs within decentralized settings while adhering to storage and privacy constraints. Our work begins with a comprehensive empirical analysis of FCGL, assessing its data characteristics, feasibility, and effectiveness, and reveals two principal challenges: local graph forgetting (LGF), where local GNNs forget prior knowledge when adapting to new tasks, and global expertise conflict (GEC), where the global GNN exhibits sub-optimal performance in both adapting to new tasks and retaining old ones, arising from inconsistent client expertise during server-side parameter aggregation. To tackle these, we propose the POWER framework, which mitigates LGF by preserving and replaying experience nodes with maximum local-global coverage at each client and addresses GEC by using a pseudo prototype reconstruction strategy and trajectory-aware knowledge transfer at the central server. Extensive evaluations across multiple graph datasets demonstrate POWER’s superior performance over straightforward federated extensions of the centralized CGL algorithms and vision-focused federated continual learning algorithms. Our code is available at https://github.com/zyl24/FCGL_POWER.
I Introduction
Graph neural networks (GNNs) are an effective approach for leveraging the relational data in graph databases for learning tasks, achieving remarkable results across diverse real-world scenarios, such as recommendation [1, 2, 3], biomedical [4, 5, 6], and finance [7, 8, 9]. However, in the era of big data, graph structures often evolve as new entities and relationships emerge, while storing or accessing massive historical node profiles (e.g., features and labels) and topologies is impractical considering database capacity and security [10, 11, 12]. As a result, GNNs often experience unacceptable performance degradation on past tasks (i.e., graphs prior to changes) when adapting to current tasks, known as the catastrophic forgetting problem, a consequence of a weak stability-plasticity tradeoff [13, 14, 15, 16]. To tackle this issue, inspired by the success of continual learning in vision tasks [17, 18, 19], various continual graph learning (CGL) studies have been proposed in recent years, all of which have demonstrated satisfactory performance [20, 21, 22, 23, 24].
Figure 1: Illustration of our FCGL paradigm, which achieves collaborative CGL through multi-round GNN parameter communications between clients and server. Each client’s evolving graph is represented by two tasks, with node colors indicating different categories.
I-A Overlooked Potential: CGL with Collective Intelligence
Despite their effectiveness, current CGL methods are based on the assumption of centralized data storage, where a single institution collects and manages the evolving graph. However, this assumption does not always hold in real-world applications. In many practical scenarios, graph-structured data is distributed across multiple decentralized sources, each capturing and responding to local changes in nodes and topology over time.
Motivating Scenario. In the e-commerce domain, various platforms maintain private evolving product networks, where nodes represent products and edges capture association relationships (e.g., co-purchases or shared reviews), both evolving as new products or associations are introduced.
In such decentralized settings, existing CGL methods face inherent limitations, as each institution can only train its GNN independently, relying solely on knowledge from its own evolving graph. This isolated learning contrasts with how humans learn, as humans benefit from collective intelligence by leveraging knowledge within groups (e.g., social media groups, academic seminars) [25]. Obviously, if an institution could leverage knowledge shared by others, it would be better positioned to mitigate catastrophic forgetting and enhance task performance. However, constraints like data privacy and commercial competition make directly aggregating multiple evolving graphs unfeasible, presenting significant challenges to achieving effective CGL with collective intelligence.
I-B First Exploration: Federated Continual Graph Learning
Building on the success of federated graph learning (FGL) [26], we introduce federated continual graph learning (FCGL) as the first practical framework for collaborative CGL in decentralized environments that adheres to storage and privacy limitations. As shown in Fig. 1, FCGL employs a multi-round communication process. The clients first train local GNNs on their private evolving graphs, and a central server then aggregates these models to form a global GNN, which has improved performance across diverse local graphs. The global model is subsequently broadcast to the clients, enabling the leveraging of collective intelligence without directly sharing massive and sensitive evolving graphs. Notably, although various federated continual learning methods [27, 28, 29, 30] have demonstrated effectiveness in vision tasks, their algorithms rely on augmentation strategies and network architectures tailored specifically for image data, rendering them inapplicable to graph data. Furthermore, for FCGL, the characteristics of multi-client evolving graphs, feasibility, and potential challenges impacting the GNN performance remain unexplored.
I-C Our Efforts: Establishing an Effective FCGL Paradigm
In order to establish an effective FCGL paradigm, we have undertaken the following efforts and contributions.
E1: In-depth Empirical Investigation of FCGL. We conduct in-depth empirical investigations of FCGL (Sec. III). From the Data perspective, we explore the data distribution of multi-client evolving graphs and identify the phenomenon of divergent graph evolution trajectories, which leads each client to develop expertise in specific classes at different stages, forming the foundation of their collective intelligence. From the Feasibility perspective, we observe that even the simplest implementation of FCGL outperforms isolated CGL methods in decentralized scenarios, demonstrating its ability to harness collective knowledge from multi-client evolving graphs. From the Effectiveness perspective, we identify two non-trivial challenges impacting GNN performance: Local Graph Forgetting (LGF), where local GNNs forget knowledge from previous tasks when adapting to new ones, and Global Expertise Conflict (GEC), arising from the multi-client graph knowledge conflicts, during the parameter aggregation across clients with divergent graph evolution trajectories. This results in sub-optimal global GNN performance on both past and current tasks.
E2: Novel and Effective FCGL Framework. Building on the insights from E1, we propose the first FCGL framework, named POWER (graPh evOlution Trajectory-aware knoWledge TransfER). Specifically, POWER is designed with two primary objectives: (1) Addressing LGF at local clients: POWER selects experience nodes from past tasks with maximum local-global coverage and replays them during local training for future tasks; and (2) Tackling GEC at central server: POWER introduces a novel pseudo prototype reconstruction mechanism, enabling the server to capture multi-client graph evolution trajectories. Leveraging this, a trajectory-aware knowledge transfer process is applied to restore the lost knowledge of global GNN during parameter aggregation.
Our Contributions: (1) Problem Identification. To the best of our knowledge, this is the first exploration of the FCGL paradigm, offering a practical solution for continual graph learning with collective intelligence under decentralized graph data management. (2) In-depth Investigation. (Sec. III) We conduct in-depth empirical investigations of FCGL from the perspectives of Data, Feasibility and Effectiveness, providing valuable insights for its development. (3) Novel Framework. (Sec. IV) We present POWER, an effective FCGL framework that tackles LGF by replaying experience nodes from past tasks with maximum local-global coverage, and handles GEC with an innovative pseudo prototype reconstruction mechanism and graph evolution trajectory-aware knowledge transfer process.
(4) State-of-the-art Performance. (Sec. V) Extensive experiments on eight datasets demonstrate that POWER consistently outperforms the federated extension of centralized CGL algorithms and vision-based federated continual learning algorithms.
II Preliminaries and Related Works
In this section, we first introduce static graphs and GNNs from a data flow perspective. We then formalize the problem definition for FCGL. Finally, we briefly review two pertinent research domains, including CGL and FGL.
II-A Notations and Problem Formalization
Static Graph and GNNs. Consider an undirected static graph $G=(\mathcal{V},\mathcal{E})$ with $\mathcal{V}$ as the node-set and $\mathcal{E}$ as the edge-set. Each node $v_{i}\in\mathcal{V}$ has an attribute vector $\mathbf{x}_{i}\in\mathbb{R}^{F}$ and a one-hot label $\mathbf{y}_{i}\in\mathbb{R}^{C}$, where $F$ is the dimension of feature and $C$ is the number of classes. The adjacency matrix (including self-loops) is $\mathbf{A}\in\mathbb{R}^{N\times N}$, where $N=|\mathcal{V}|$ is the number of nodes. $\mathcal{N}(v_{i})$ represents the neighbor nodes set of node $v_{i}$. Most GNNs can be subsumed into a framework named Message Passing Neural Network (MPNN) [31], which obtains the representation of a specific node by recursively aggregating and transforming the representations of its neighbors. Specifically, the representation of node $v_{i}$ at the $l$-th layer is denoted as $\mathbf{h}_{i}^{l+1}$, which can be formalized as follows:
$\mathbf{h}_{i}^{l+1}=\text{UPD}(\mathbf{h}_{i}^{l},\text{AGG}(\{\mathbf{h}_{j}^{l}:v_{j}\in\mathcal{N}(v_{i})\})),$ (1)
where $h_{i}^{0}=x_{i}$, $h_{i}^{l}$ is the representation of node $v_{i}$ in the $l$-th (previous) layer, $\text{AGG}(\cdot)$ aggregates the neighbor representations, and $\text{UPD}(\cdot,\cdot)$ updates the representation of the current node using its representation and the aggregated neighbor representation at the previous layer. The semi-supervised node classification task is based on the topology of labeled set $\mathcal{V}_{L}$ and unlabeled set $\mathcal{V}_{U}$, and the nodes in $\mathcal{V}_{U}$ are predicted based on the model supervised by $\mathcal{V}_{L}$.
Problem Formalization of FCGL. For FCGL, there is a trusted central server and $K$ clients. Specifically, for the $k$-th client, we use $\mathcal{T}^{k}=\{T^{k}_{1},T^{k}_{2},...,T^{k}_{M}\}$ to denotes the set of its tasks, where $M$ is the number of tasks. Then, the local private evolving graph with $M$ tasks can be formalized as follows:
$\mathcal{G}^{k}=\{G^{k}_{1},G^{k}_{2},...,G^{k}_{M}\};$ (2)
$G^{k}_{t}=G^{k}_{t-1}+\Delta G^{k}_{t},$
where $G^{k}_{t}=(\mathcal{V}^{k}_{t},\mathcal{E}^{k}_{t})$ is the graph given in task $T^{k}$, and $\Delta G^{k}_{t}=(\Delta\mathcal{V}^{k}_{t},\Delta\mathcal{E}^{k}_{t})$ is the change of the node and edge set in task $T^{k}_{t}$. The purpose of graph continual learning is to train models $\{\text{GNN}_{\mathbf{\Theta}_{1}^{k}},\text{GNN}_{\mathbf{\Theta}_{2}^{k}},\dots,\text{GNN}_{\mathbf{\Theta}_{M}^{k}}\}$ from streaming data where $\mathbf{\Theta}^{k}_{t}$ is the parameter of the GNN model in task $T^{k}_{t}$.
In this paper, we focus on the semi-supervised node classification task under the class-incremental setting, where unseen classes are continually introduced as the graph evolves [32]. To simplify the problem, we assume that all clients have an equal number of tasks $M$, with each task introducing $c$ new node categories and undergoing $R$ rounds of FCGL communication. Thus, each client will have encountered $C\!=\!c\!\times\!M$ categories after training on the last task ($t\!=\!M$). Consider the widely-used FedAvg [33] aggregation strategy in federated learning for vision tasks. Adapting this strategy within the FCGL framework, the training process unfolds as follows:
  1. 1.
    Initialization. This step occurs only at the first communication round ($r\!=\!1$) of the first task ($t\!=\!1$). The central server sets the local GNN parameters of $K$ clients to the global GNN parameters $\mathbf{\Theta}^{\text{g}}$, using: $\mathbf{\Theta}^{k}\leftarrow\mathbf{\Theta}^{\text{g}}\,\forall k$.
  2. 2.
    Local Updates. Each local GNN performs training on the current local data $G_{t}^{k}$ to minimize the task loss $\mathcal{L}(G_{t}^{k};\mathbf{\Theta}_{t}^{k})$, and then updating the parameters: $\mathbf{\Theta}^{k}\leftarrow\mathbf{\Theta}^{k}-\eta\nabla\mathcal{L}$.
  3. 3.
    Global Aggregation. After local training, the server aggregates the local knowledge with respect to the number of training instances, i.e., $\mathbf{\Theta}^{\text{g}}\leftarrow\frac{N_{k}}{N}\sum_{k=1}^{K}\mathbf{\Theta}^{k}$ with $N=\sum_{k}N_{k}$, and distributes the global parameters $\mathbf{\Theta}^{\text{g}}$ to local clients selected at the next round.
For each task $t$, this process iterates between steps 2 and 3 until reaching the final round $R$. This iterative cycle continues until the completion of the last round ($r\!=\!R$) of the last task ($t\!=\!M$), facilitating collaborative CGL by parameter sharing without the exchange of private local data.
II-B Continual Graph Learning
CGL introduces a novel paradigm where GNNs learn from dynamically evolving graphs while retaining prior knowledge. The primary challenge in CGL, catastrophic forgetting, manifests as significant performance degradation on historical tasks as the GNN adapts to new data. Mainstream CGL research is categorized into three primary approaches: (1) Experience Replay focuses on selecting and storing a minimal set of experiential data from previous tasks. Its core objective is to devise strategies that efficiently capture and retain representative data within constrained memory limits for future task training. This involves deciding which data to replay and how to utilize these samples effectively for subsequent re-execution. Pioneering efforts in this field include ERGNN [20], which stores experience data at the node level, devises node selection strategies such as coverage and social influence maximization and incorporates node classification losses during new task training. DSLR [23] enhances node selection with a coverage-based diversity strategy for class representativeness and intra-class diversity and refines node topology through graph structure learning. Additionally, recent studies [34, 35, 24] propose the experience data selection strategy at the subgraph or ego-network level, which explicitly preserves topological structure information for replay and achieves satisfactory performance. (2) Parameter Regularization alleviates catastrophic forgetting by introducing a regularization term during the training of new tasks to preserve the parameters crucial for previous tasks. For instance, TWP [21] introduces a novel regularization term that preserves critical GNN parameters for node semantics and topology aggregation, considering both node-level and topology-level knowledge from previous tasks. Moreover, SSRM [22] introduces a novel regularization-based technique to mitigate the catastrophic forgetting caused by structural shifts. (3) Architecture Isolation assigns a dedicated set of GNN parameters to each task, ensuring that learning a new task does not interfere with the parameters of the old task. Studies in this line  [36, 37] effectively acquire knowledge from new graphs by expanding the GNN architecture when its capacity is insufficient to meet the design requirements.
II-C Federated Graph Learning
Motivated by the success of federated learning in computer vision and natural language processing [38] with the demands for distributed graph learning in the real world, FGL has received growing attention from researchers. From the data and task perspectives, mainstream FGL studies can be divided into two settings: (1) Graph-FL: In this setting, each client independently collects multiple graphs to address graph-level downstream tasks, such as graph classification collaboratively. The main challenge of Graph-FL is to avoid information interference between graph datasets from different clients, particularly when handling multi-domain data. For example, GCFL+[39] introduces a GNN gradient pattern-aware technique for dynamic client clustering, which mitigates knowledge conflicts caused by structural and feature heterogeneity across clients’ graph data. (2) Subgraph-FL: In this setting, each client holds a subgraph of an implicit global graph, aiming to solve node-level tasks such as node classification. The main challenges of Subgraph-FL include subgraph heterogeneity (i.e., performance degradation due to node and topology variations across subgraphs) and missing edges (i.e., lost connections between subgraphs due to distributed storage). To address subgraph heterogeneity, Fed-PUB [40] improves local GNNs by computing similarities from random graph embeddings for weighted server aggregation, with clients using personalized sparse masks to selectively update subgraph parameters. FedGTA [41] encodes topology into smoothing confidence and graph moments to improve model aggregation. Moreover, other studies [42, 43, 44, 45] also achieve satisfactory results on this challenge. To address missing edges, FedSage+[26] integrates node representations, topology, and labels across subgraphs while training a neighbor generator to restore missing links, achieving subgraph-FL with robust and complete graph information. Several other works [46, 47, 48] also show strong performance in addressing this challenge. Further detailed insights into FGL research can be found in comprehensive surveys [49, 50, 51] and extensive benchmark studies [52, 53, 54].
Figure 2: Experimental results of our empirical study. (a) Node label distribution of each client’s class-incremental tasks, each exhibiting divergent evolution trajectories. (b) Comparative analysis of three CGL methods in isolated versus federated settings, presenting AM and FM metrics in the upper and lower parts, respectively. (c) Performance variation of Client 1’s local GNN (Single GNN) during training on Client 1 Task 3, with markers denoting receiving global parameters from the server. (d) Performance comparison between clients and the server (Multiple GNNs) during training on Client 1 Task 3, indicating that the global GNN can be improved by local GNNs with expertise.
III Empirical Investigation
In this section, we present a thorough empirical investigation of the proposed FCGL paradigm, structured around three key questions from different perspectives. From the Data perspective, to better understand the complexities of graph data in FCGL contexts, we answer Q1: What patterns and trends emerge in data distribution across multi-client evolving graphs in FCGL settings? From the Feasibility perspective, to demonstrate the advantages of FCGL using collective intelligence, we answer Q2: Compared to isolated CGL, does a naive FCGL approach (i.e., CGL integrated with FedAvg) achieves better performance for decentralized evolving graphs? From the Effectiveness perspective, to reveal the bottleneck of naive FCGL and realize efficient FCGL paradigm, we answer Q3: What non-trivial challenges limit the performance of GNNs within the FCGL framework? Details for the described experimental setup, algorithms, and metrics refer to Sec. V-A.
To address Q1, we simulate multiple decentralized evolving graphs using a two-step approach. First, we partition the Cora dataset [55] into three subgraphs using the Louvain algorithm [56], which is widely utilized in various FGL studies [41, 26, 48]. These subgraphs are distributed among three different clients. Second, following the standard experimental settings of CGL [21, 20, 23], each client’s local subgraph is further divided into three class-incremental tasks. Each task comprises two classes of nodes, discarding any surplus classes and removing inter-task edges while preserving only intra-task connections. For each client, we present its tasks with node label distributions in Fig. 2 (a).
Observation 1. Although all these clients eventually learn across six classes, they follow divergent graph evolution trajectories, varying both in sample quantity (e.g., Clients 1 and 2) and in the sequence of class learning (e.g., Clients 1 and 3). Intuitively, these clients tend to develop expertise in specific classes at different stages, forming the basis of the collective intelligence harnessed by the FCGL paradigm.
To address Q2, we assess the performance of three CGL algorithms, including ERGNN [20], TWP [21], and DSLR [23], under two scenarios: isolated training (referred to as Iso-ERGNN, Iso-TWP, and Iso-DSLR) and federated training using the FedAvg [33] algorithm (referred to as Fed-ERGNN, Fed-TWP, and Fed-DSLR). In the isolated setting, no client-server communication occurs, while the federated variants implement naive FCGL strategies that simply aggregate parameters from local GNNs. We assessed their performance using two metrics: the accuracy mean (AM) and the forgetting mean (FM), where higher AM values denote better adaptation to new tasks, and lower FM values denote less forgetting of previously learned tasks. The training protocol for each task includes 10 communication rounds, with 3 local epochs per round. As depicted in Fig. 2 (b), the federated variants consistently outperformed their isolated counterparts in both AM and FM metrics. This enhanced performance is attributed to FCGL’s capability to leverage diverse knowledge from evolving graph data across clients, thereby reducing task forgetting through effective cross-client knowledge transfer (e.g., alleviating the knowledge forgetting of Client 1 Task 1 and Client 2 Task 1 through insights gained from Client 3 Task 3).
Observation 2. Even a basic implementation of FCGL can substantially enhance the performance of existing CGL algorithms on decentralized evolving graphs.
Figure 3: Overview of our proposed POWER framework. We follow a class-incremental setting, where each client collects an evolving graph into which unseen classes are continually introduced. Node colors indicate different labels. (1) To tackle LGF, POWER stores experience nodes with maximum local-global coverage in the last round of the old task and replays them in each round of the new task; (2) To tackle GEC, POWER reconstructs pseudo prototypes via client-server collaboration in the first round of each task. Based on this, POWER constructs a global buffer graph and applies trajectory-aware knowledge transfer to recover the global GNN’s lost knowledge in each round of each task.
To address Q3, we evaluate the naive FCGL algorithm’s performance across Client 1’s three tasks during the training of Task 3. For readability, we present the results obtained from Fed-ERGNN, as similar trends are also shown in Fed-TWP and Fed-DSLR. Our analysis is further concentrated on rounds 4 and 5. From the Single GNN aspect, we analyze performance variations in Client 1’s local GNN, particularly during the critical phases of global parameter reception and local training, motivated by findings in existing literature [28, 21], which suggest that shifts in parameter values play a significant role in causing knowledge forgetting. This perspective enables a closer examination of how individual GNNs react to external updates. As shown in Fig. 2 (c), each global parameter reception phase is marked with scatter points, while lines represent the 3-epoch local training phases. From the Multiple GNNs aspect, we compare the performance of all local GNNs and the aggregated global GNN. The results are depicted in Fig. 2 (d).
Observation 3. Building on these two aspects, we reveal the following two Non-trivial Challenges limit the performance of GNNs within the FCGL framework:
  1. 1.
    Local Graph Forgetting (LGF). From the Single GNN aspect, local training on the current task (i.e., Client 1 Task 3) significantly impairs performance on earlier tasks (i.e., Client 1 Tasks 1 and 2). This demonstrates that in the FCGL scenario, the local GNN experiences local graph forgetting similar to that of centralized CGL.
  2. 2.
    Global Expertise Conflict (GEC). From the Single GNN aspect, incorporating global parameters can degrade the performance on the current task (i.e., Client 1 Task 3) while improving the performance on previous tasks (i.e., Client 1 Tasks 1 and 2). However, these improvements are limited when considering Multiple GNNs aspect. For example, in Task 1, the global GNN is outperformed by the local GNN of Client 3; in Task 2, it is outperformed by the local GNN of Client 2. This phenomenon occurs because each client’s local GNN develops expertise shaped by its unique evolution trajectory, and this expertise leads to conflicts during parameter aggregation, where the global GNN fails to capture this expertise adequately. Thus, GEC hinders FCGL’s ability to adapt to new tasks and mitigate forgetting of previous ones.
In summary, both LGF and GEC present non-trivial challenges that critically restrict the performance of GNNs in FCGL. A meticulous design for tackling these challenges is imperative to realizing an effective FCGL framework.
IV POWER: The Proposed Framework
In this section, we introduce POWER, the first FCGL framework designed for learning on multiple decentralized evolving graphs. We first provide a structured overview of POWER in Fig. 3. Afterward, we delineate the architecture of POWER based on its two design objectives. Specifically, in Sec. IV-A, we detail the maximum local-global coverage-based experience node selection and replay mechanisms that POWER employs to address LGF at local clients. In Sec. IV-B, we expound on the pseudo prototype reconstruction mechanism and the graph evolution trajectory-aware knowledge transfer process, which is employed to tackle GEC at central server.
IV-A Addressing Local Graph Forgetting
To tackle the challenge of LGF, POWER employs a strategic selection and storage of experience nodes from prior tasks, enabling targeted replay during local training on new tasks. We begin by introducing the experience node selection strategy.
Experience Node Selection. To ensure that selected experience nodes capture knowledge from previous tasks, we build upon the Coverage Maximization (CM) strategy [20] with further modifications tailored to FCGL settings. The conventional CM is based on the intuitive assumption that nodes within the same category share similar properties in the embedding space, with many nodes of the same class naturally gathering near their class’s representative nodes. As for FCGL, we can utilize the embeddings from both the local and global GNNs to fully consider both local and global knowledge. Specifically, consider the $k$-th client training on its $t$-th local task $T_{t}^{k}$; the local graph is denoted as $G_{t}^{k}$. At the last training round, we calculate the local-global node embeddings $\mathbf{Z}$ as follows:
$\mathbf{H}=\text{GNN}($ $\mathbf{X}_{t}^{k},\mathbf{A}_{t}^{k}|\mathbf{\Theta}^{k}),$ (3)
$\mathbf{H}^{\text{g}}=\text{GNN}$ $(\mathbf{X}_{t}^{k},\mathbf{A}_{t}^{k}|\mathbf{\Theta}^{\text{g}}),$
$\mathbf{Z}=\alpha\mathbf{H}+$ $(1-\alpha)\mathbf{H}^{\text{g}},$
where $\mathbf{H}$ and $\mathbf{H}^{\text{g}}$ are node embeddings from local and global GNNs with parameters $\mathbf{\Theta}^{k}$ and $\mathbf{\Theta}^{\text{g}}$, respectively. $\alpha$ serves as a trade-off factor for combining local and global embeddings. The local-global coverage for each labeled node $v_{i}$ in the training set $\mathcal{V}^{\text{lbl}}$ is denoted as $C(v_{i})$, defined as:
$C(v_{i})=|\{v_{j}\mid v_{j}\in\mathcal{V}^{\text{lbl}},\mathbf{y}_{i}=\mathbf{y}_{j},d(\mathbf{z}_{i},\mathbf{z}_{j})<\epsilon E(v_{i})\}|,$ (4)
where $d(\cdot,\cdot)$ measures the Euclidean distance between the embeddings of $v_{i}$ and $v_{j}$, $\epsilon$ serves as a threshold distance, and $E(v_{i})$ denotes the average pairwise distance between embeddings of training nodes in the same class as $v_{i}$.
Subsequently, for each class $c$ in the current task, we select $b$ experience nodes from the set of nodes in class $c$ with maximum local-global coverage. The experience node set of class $c$ is denoted as $\mathcal{B}_{c}$, which is formulated as follows:
$\mathcal{B}_{c}=\operatorname*{argmax}\nolimits_{\{v_{c_{1}},...,v_{c_{b}}|v_{c_{1}},...,v_{c_{b}}\in\mathcal{V}_{c}^{\text{lbl}}\}}\sum\nolimits_{i=1}^{b}C(v_{c_{i}}),$ (5)
where $\mathcal{V}_{c}^{\text{lbl}}=\{v_{i}|v_{i}\in\mathcal{V}^{\text{lbl}},\mathbf{y}_{i}\!=\!c\}$, $b$ represents the buffer size allocated per task and class, constrained by the storage capacity budget. In our experiments, we set $b=1$ as the default. Eq. 5 poses an NP-hard challenge, which we address using a greedy algorithm. At each iteration, the algorithm picks the node with the highest local-global coverage from the remaining unselected nodes in $\mathcal{V}_{c}^{\text{lbl}}$, repeating until $b$ nodes are selected.
Finally, for each class $c$ in task $T_{t}^{k}$, the experience node set $\mathcal{B}_{c}$ is stored into the local buffer $\mathcal{B}$ for replay (i.e., $\mathcal{B}=\mathcal{B}\cup\mathcal{B}_{c}$).
Experience Node Replay. POWER’s local training process pursues two primary objectives. First, to adapt to the current task $T_{t}^{k}$, the local GNN minimizes the cross-entropy loss over the labeled training set $\mathcal{V}^{\text{lbl}}$, as formulated below:
[…]
$\mathcal{L}_{\text{new}}=-\!\!\sum\nolimits_{v_{i}\in\mathcal{V}^{\text{lbl}}}\mathbf{y}_{i}\log$ $\hat{\mathbf{y}}_{i}+(1-\mathbf{y}_{i})\log(1-\hat{\mathbf{y}}_{i}),$
where $\hat{\mathbf{Y}}$ denotes the soft labels predicted by the local GNN for the current task nodes, $\mathbf{\Theta}^{k}$ represents the local GNN parameters, and $\mathbf{y}_{i}$ is the ground-truth label for a specific training node $v_{i}$.
When the current task is not the first task (i.e., $t\neq 1$), the client also optimizes a second objective, aiming at retaining knowledge from previously learned tasks. To achieve this, the client replays experience nodes stored in the local buffer $\mathcal{B}$, which captures representative samples from prior tasks. The replay objective is the cross-entropy loss over these stored nodes, which can be calculated as follows:
$\hat{\mathbf{Y}}=\text{GNN-CLS}(\mathbf{X}_{\mathcal{B}},\mathbf{I}~{}|~{}\mathbf{\Theta}^{k}),$ (7)
$\mathcal{L}_{\text{old}}=-$ $\sum\nolimits_{v_{i}\in\mathcal{B}}\mathbf{y}_{i}\log\hat{\mathbf{y}}_{i}+(1-\mathbf{y}_{i})\log(1-\hat{\mathbf{y}}_{i}),$
where $\hat{\mathbf{Y}}$ represents the soft labels predicted by the local GNN for the buffer nodes, while $\mathbf{X}_{\mathcal{B}}$ denotes the feature matrix of nodes in the local buffer, and $\mathbf{y}_{i}$ is the ground truth for a specific buffer node $v_{i}$. Notably, since we store only the node features without including any topological structures, the adjacency matrix for buffer nodes is set as the identity matrix $\mathbf{I}$, meaning that each buffer node is predicted in isolation.
Finally, the loss function of local training is defined as:
$\mathcal{L}=\beta\mathcal{L}_{\text{new}}+(1-\beta)\mathcal{L}_{\text{old}},$ (8)
where $\beta$ denotes the trade-off parameter to flexibly balance adaptation to new tasks with a replay of previous tasks.
IV-B Tackling Global Expertise Conflict
As discussed in Sec. III, GEC arises from knowledge conflicts during the aggregation of parameters from local GNNs trained on divergent graph evolution trajectories. Consequently, for certain classes, the global model exhibits sub-optimal performance compared to clients with expertise in those classes. Motivated by this, POWER employs two modules to address GEC: (1) Pseudo Prototype Reconstruction, which enables the server to discern each client’s evolution trajectory and expertise, and (2) Trajectory-aware Knowledge Transfer, which utilizes multi-client trajectories and expertise to recover the lost knowledge of the global GNN.
Pseudo Prototype Reconstruction Mechanism. First, to capture the expertise of each client, POWER calculates the prototype (i.e., averaged node feature) for each class on the client side. Specifically, consider the $k$-th client on its first communication round of $t$-th local task $T_{t}^{k}$. We denote the labeled node set as $\mathcal{V}^{\text{lbl}}$, with $\mathcal{V}^{\text{lbl}}_{c}=\{v_{i}\mid v_{i}\in\mathcal{V}^{\text{lbl}},y_{i}=c\}$ representing the subset of nodes labeled as class $c$. The prototype of class $c$ is denoted as $P_{c}$, which is defined as:
$P_{c}=\frac{1}{|\mathcal{V}^{\text{lbl}}_{c}|}\sum\nolimits_{v_{i}\in\mathcal{V}^{\text{lbl}}_{c}}\mathbf{x}_{i},$ (9)
However, directly transmitting class prototypes to the server incurs substantial privacy risks. To mitigate this, POWER adopts a privacy-preserving strategy wherein clients send prototype gradients instead, allowing the server to reconstruct pseudo prototypes via gradient matching. This approach enhances security for two key reasons: (1) the prototype gradients are derived from a randomly initialized network, making it challenging to extract meaningful information from gradients alone without additional context, and (2) the pseudo prototypes produced on the server only mimic the gradient patterns of the true prototypes, lacking identifiable numerical features that could be recognized by humans. Specifically, the client feeds the computed prototype $\mathbf{P}_{c}$ into an $L$-layer randomly initialized gradient encoding network $\Gamma$ with parameters $\{\mathbf{\Omega}_{i}\}_{i=1}^{L}$ to obtain predictions. We then backpropagate the cross-entropy loss to compute the gradient $\nabla\Gamma_{c}$, whose $i$-th element $\nabla_{\mathbf{\Omega}_{i}}$ is formulated as:
$\nabla_{\mathbf{\Omega}_{i}}\Gamma_{c}$ $=\!\nabla_{\mathbf{\Omega}_{i}}(-\mathbf{y}_{c}\log\hat{\mathbf{y}}_{c}\!+\!(1\!-\!\mathbf{y}_{c})\log(1\!-\!\hat{\mathbf{y}}_{c})),$ (10)
where $\hat{\mathbf{y}}_{c}=\Gamma(P_{c}|\{\mathbf{\Omega}_{i}\}_{i=1}^{L})$ denotes the predicted soft label and $\mathbf{y}_{c}$ is the one-hot ground-truth vector representing class $c$.
When the server receives the prototype gradient $\nabla\Gamma$, it reconstructs pseudo prototypes via gradient matching. Specifically, consider a trainable pseudo prototype $\hat{\mathbf{P}}$ initialized by a standard Gaussian noise $\mathcal{N}(0,1)$, the server obtains the pseudo prototype prediction by feeding $\hat{\mathbf{P}}$ into a prototype reconstruction network $\Gamma$, which is the same as the gradient encoding network used by all local clients (i.e., also parameterized by $\{\mathbf{\Omega}_{i}\}_{i=1}^{L}$). Subsequently, the server backpropagates the cross-entropy loss to compute the pseudo prototype gradients $\nabla\hat{\Gamma}$, whose $i$-th element $\nabla_{\mathbf{\Omega}_{i}}$ is formulated as follows:
$\nabla_{\mathbf{\Omega}_{i}}\hat{\Gamma}$ $=\!\nabla_{\mathbf{\Omega}_{i}}(-\mathbf{y}\log\hat{\mathbf{y}}\!+\!(1\!-\!\mathbf{y})\log(1\!-\!\hat{\mathbf{y}})),$ (11)
where $\hat{\mathbf{y}}=\Gamma(P|\{\mathbf{\Omega}_{i}\}_{i=1}^{L})$ is the predicted soft label of pseudo prototype and $\mathbf{y}$ is the one-hot vector of a specific class $c$. Notably, the class $c$ of prototype gradient $\nabla\Gamma$ can be inferred with the symbol of prototype gradient (i.e., $c=\text{argmin}_{i}\nabla_{\Omega_{L_{i}}}\Gamma$).
Based on Eqs. (10) and (11), the trainable pseudo prototype is optimized via the gradient matching loss $\mathcal{L}_{\text{gm}}$, defined as:
$\mathcal{L}_{\text{gm}}=\sum\nolimits_{i=1}^{L}|\!|\nabla_{\mathbf{\Omega}_{i}}\Gamma-\nabla_{\mathbf{\Omega}_{i}}\hat{\Gamma}|\!|^{2}.$ (12)
Finally, the reconstructed pseudo prototypes are stored in global buffer $\mathcal{B}^{\text{g}}$ (i.e., $\mathcal{B}^{\text{g}}=\mathcal{B}^{\text{g}}\cup\{\hat{\mathbf{P}}\}$).
Second, we define the evolution trajectory of each client through a cumulative label distribution, which aggregates task-specific label distributions across the client’s task sequence, applying a gradual decay to earlier tasks to simulate LGF during local training. Formally, the evolution trajectory for the $k$-th client on task $T^{t}$ is represented as $\mathbf{q}_{t}^{k}$, defined as follows:
$\forall i\in\{1,...,t\},\mathbf{p}_{i}^{k}$$\propto\{\sum\nolimits_{v_{j}\in\mathcal{V}^{\text{lbl}}}\mathbf{1}_{y_{j}=c}\}_{c\in\mathcal{C}},$ (13)
$\mathbf{q}_{t}^{k}=$ $\sum\nolimits_{i=1}^{t}\phi^{t-i}\mathbf{p}_{i}^{k},$
where $\phi$ represents the decay coefficient that attenuates the contributions of past tasks, $\mathcal{C}$ denotes the total classes, $\mathbf{q}_{i}^{k}$ denotes the label distribution of task $T_{i}^{k}$, and $\mathbf{1}_{y_{j}=c}$ denotes an indicator expression, taking the value 1 if $y_{j}=c$ holds, and 00 otherwise. Notably, Eq. (13) can be computed iteratively without storing any nodes from previous tasks.
Trajectory-aware Knowledge Transfer. Our key insight is to utilize multi-client expertise and trajectories to recover the lost knowledge of the global GNN. First, based on the global buffer $\mathcal{B}^{\text{g}}$, the server constructs a global buffer graph $G^{\text{g}}$ via the K-Nearest Neighbor strategy to explore the global input space. Specifically, the feature matrix of the global buffer is denoted as $\mathbf{X}^{\text{g}}$, the constructed adjacency $\mathbf{A}^{\text{g}}$ can be calculated as:
$\mathbf{H}\!=\!\sigma(\mathbf{X}^{\text{g}}\mathbf{X}^{\text{g}^{T}}\!),$ (14)
$\mathbf{A}^{\text{g}}[u,v]$ ={$&1,\text{if }v\in\text{TOPK}(\bar{k},\mathbf{H}[u,:]),\\ &0,\text{otherwise},\end{aligned}\right.$
where $\sigma(\cdot)$ represents the element-wise sigmoid function, and $\text{TOPK}(\bar{k},\cdot)$ selects the index of the first $\bar{k}$ largest items.
Subsequently, POWER leverages the global buffer graph $\mathcal{G}^{\text{g}}$ along with the set of clients’ evolution trajectories $\mathbf{q}_{t}=\{\mathbf{q}_{t}^{k}\}_{k=1}^{K}$ to facilitate knowledge transfer. The extent to the
1
Input: Number of tasks $T$, each task communication rounds $R$, local epochs $E$, global epochs $E_{g}$, global GNN parameter $\mathbf{\Theta}^{\text{g}}$
Output: Local GNN parameters $\{\mathbf{\Theta}^{k}\}_{k=1}^{K}$
2
3for each training task $t$ in $1,...,T$ do
4       for each communication round $r$ in $1,...,R$ do
5             for each client $k=1,...,K$ do
6                   $\mathbf{\Theta}^{k}$, $\nabla\Gamma_{t}^{k}$, $\mathbf{q}_{t}^{k}\leftarrow$C_Exec($k,t,r,\mathbf{\Theta}^{\text{g}}$)
7             end for
            $\mathbf{\Theta}^{\text{g}}\leftarrow$S_Exec($t$, $r$, $\{\nabla\Gamma_{t}^{k}\}_{k=1}^{K}$, $\{\mathbf{q}_{t}^{k}\}_{k=1}^{K}$, $\{\mathbf{\Theta}^{k}\}_{k=1}^{K}$)
              // Receive uploaded messages
8
9       end for
10
11 end for
12
/* Client Execution */
13 Function C_Exec($k,t,r,\mathbf{\Theta}^{\text{g}}$):
       $\mathbf{\Theta}^{\text{k}}\leftarrow\mathbf{\Theta}^{\text{g}}$
        // Update local GNN parameters
14       for each local training epoch $e$ in $1,...,E$ do
             Perform local training and replaying.
              // Eq. 8
15
16       end for
17      if $r=R$ then
18             for each class $c$ in task $T_{t}^{k}$ do
                   Select experience nodes $\mathcal{B}_{c}$
                    // Eqs.3, 4, 5
19
                  $\mathcal{B}\leftarrow\mathcal{B}\cup\mathcal{B}_{c}$
                    // Store to local buffer
20
21             end for
22
23       end if
24      if $r=1$ then
25             $\nabla\Gamma_{t}^{k}\leftarrow\{\}$
26            for each class $c$ in task $T_{t}^{k}$ do
                   Compute prototype gradients $\nabla\Gamma_{c}$
                    // Eq.10
27                   $\nabla\Gamma_{t}^{k}\leftarrow\nabla\Gamma_{k}^{t}\cup\nabla\Gamma_{c}$
28             end for
29
            Compute evolution trajectory $\mathbf{q}_{t}^{k}$
              // Eq. 13
30             return $\mathbf{\Theta}^{k}$, $\nabla\Gamma_{t}^{k}$, $\mathbf{q}_{t}^{k}$
31      else
32             return $\mathbf{\Theta}^{k}$, None, None
33       end if
34
35 end
36
/* Server Execution */
37
38Function S_Exec($t$, $r$, $\nabla\Gamma_{t}$, $\mathbf{q}_{t}$, $\mathbf{\Theta}$):
39       if $r=1$ then
40             for $\nabla\Gamma_{t}^{k}$ in $\nabla\Gamma_{t}$ do
41                   for $\nabla\Gamma_{t_{c}}^{k}$ in $\nabla\Gamma_{t}^{k}$ do
                         Reconstruct $\hat{\mathbf{P}}$ via $\nabla\Gamma_{t_{c}}^{k}$
                          // Eqs.11,12
42
                        /* Store to global buffer */
43                         $\mathcal{B}^{\text{g}}\leftarrow\mathcal{B}^{\text{g}}\cup\{\hat{\mathbf{P}}\}$
44                   end for
45
46             end for
47
48       end if
49      for each global training epoch $e$ in $1,...,E_{g}$ do
             Construct global buffer graph $G^{\text{g}}$
              // Eq. LABEL:eq:_knn
             Perform Knowledge transfer
              // Eq. IV-B
50
51       end for
52      return Global GNN parameters $\mathbf{\Theta}^{\text{g}}$
53
54 end
55
Algorithm 1 POWER-Clients and Server Execution
global model’s integration of expertise from each local GNN is guided by the client’s evolution trajectory. Specifically, the knowledge transfer loss for training on task $T_{t}$ is defined as:
$\forall k\in\{1,..$$.,K\},\hat{\mathbf{Y}}^{k}=\text{GNN-CLS}(\mathbf{X}^{\text{g}},\mathbf{A}^{\text{g}}|\mathbf{\Theta}^{k}),$
[…]
$\mathcal{L}_{\text{div}}=\sum_{c=1}^{C}\sum_{k=1}^{K}$ $\sum_{v_{i}\in\mathcal{V}^{\text{g}}}\mathbf{1}(\mathbf{y}_{i}=c)\frac{\mathbf{q}_{t}^{k}(c)}{\sum\nolimits_{j=1}^{K}\mathbf{q}_{t}^{j}(c)}\text{KL}(\hat{\mathbf{y}}_{i}^{\text{g}}|\!|\hat{\mathbf{y}}_{i}^{k}),$
[…]
The full algorithm for POWER is outlined in Algorithm 1.
V Experiments
TABLE I: Statistical information of the experimental datasets, where “#Classes pT” denotes the number of classes per task.
Datasets #Nodes #Features #Edges #Classes #Classes pT #Tasks #Clients Train/Val/Test Description
Cora 2,708 1,433 5,429 7 2 3 3 20%/40%/40% Citation Network
CiteSeer 3,327 3,703 4,732 6 2 3 3 20%/40%/40% Citation Network
OGB-arxiv 169,343 128 231,559 40 10 4 5 60%/20%/20% Citation Network
Computers 13,381 767 245,778 10 2 5 3 20%/40%/40% Co-purchase Network
Physics 34,493 8,415 247,962 5 2 2 10 20%/40%/40% Co-authorship Network
Squirrel 5,201 2,089 216,933 5 2 2 10 48%/32%/20% Wiki-page Network
Flickr 89,250 500 899,756 7 2 3 10 60%/20%/20% Image Network
Roman-empire 22,662 300 32,927 18 6 3 5 50%/25%/25% Article Syntax Network
In this section, we present a comprehensive evaluation of POWER. We first introduce 8 benchmark graph datasets across 6 domains and the simulation strategy for the FCGL scenario. Then, we present 10 evaluated state-of-the-art baselines. Moreover, we provide detailed metrics description and experiment settings (Sec. V-A). After that, we aim to address the following questions: Q1: Compared with state-of-the-art baselines, can POWER achieve better predictive performance under FCGL scenario? (Sec. V-B) Q2: Does POWER demonstrate superiority in terms of efficiency? (Sec. V-C) Q2: How much does each module in POWER contribute to the overall performance? (Sec. V-D) Q4: Can POWER remain overall robust under changes in hyperparameters? (Sec. V-E) Q5: How does POWER perform under sparse settings like missing features/edges/labels and low client participation rate? (Sec. V-F)
V-A Experimental Setup
Datasets and Simulation Strategy. We evaluate POWER on 8 public benchmark graph datasets across 6 domains, including three citation networks (Cora, Citeseer [55] and OGB-arxiv [57]), one co-purchase network (Computers [58]), one co-authorship network (Physics [58]), one wiki-page network (Squirrel [59]), one image network (Flickr [60]), and one article syntax network (Roman-empire [61]). Our strategy for simulating multiple decentralized evolving graphs has been described in detail in Sec. III Q1. Notably, for each dataset, based on the number of nodes and categories, we define varying numbers of clients, task-specific category counts, and the proportions for the training, validation, and test sets. For more detailed statistical information, please refer to Table. I.
Baselines. We summarize 10 baselines into 3 categories as follows: (1) FL/FGL Fine-tuning, including FedAvg [33], FedSage+ [26], Fed-PUB [40] and FedGTA [41], which simply replace the local client training process of the FL/FGL algorithm with fine-tuning on sequential tasks; (2) Federated CGL, including Fed-ERGNN, FedTWP and Fed-DSLR, which combines the naive FedAvg algorithm with three representative CGL methods, ERGNN [20], TWP [21] and DSLR [23], respectively; (3) Federated Continual Learning for CV, which includes GLFC [27], TARGET [29] and LANDER [30].
Evaluation Metrics. We adopt two primary metrics commonly used in continual learning benchmarks [20, 21, 23] to evaluate model performance across sequential tasks: (1) Accuracy Mean (AM), defined as $\text{AM}=\frac{1}{T}\sum_{i=1}^{T}\mathbf{A}_{T,i},$ and (2) Forgetting Mean (FM), given by $\text{FM}=\frac{1}{T-1}\sum_{i=1}^{T-1}(A_{i,i}-A_{T,i}),$ where $T$ represents the total number of tasks. In standard CGL, $\mathbf{A}_{i,j}$ denotes the accuracy on task $j$ following the completion of task $i$. However, within our FCGL framework, each client’s local GNN independently predicts its own task, and we compute a weighted average of $\mathbf{A}_{i,j}$ across clients, proportional to their sample counts. AM measures the average accuracy achieved on each task after training on the final task, where higher values indicate stronger overall performance. FM, by contrast, captures the mean reduction in performance on each task as additional tasks are learned, with lower values suggesting greater retention.
Model Architecture. We employ a 2-layer GAT [62] with 64 hidden units as the backbone of clients and central server. Notably, model-specific baselines like FedSage+ [26] adhere to the custom architectures specified in their original papers. For the POWER framework, the gradient encoding network and prototype reconstruction network are implemented as 4-layer MLPs with 128, 128, and 64 hidden units. To ensure compatibility with graph data or matrix inputs, we replace the visual models used in the baseline comparisons with their respective equivalent layers of GAT or MLP.
Hyperparameters. In each task, we conduct 10 communication rounds, each comprising 3 local training epochs. These epochs utilize the Adam optimizer [63] with hyperparameters set to a learning rate of $1\times 10^{-2}$, weight decay of $5\times 10^{-4}$, and a dropout rate of 0.5. For all considered baselines, we adopt hyperparameters from the original publications whenever possible. In cases where these are not specified, we employ an automated hyperparameter optimization using the Optuna framework [64]. For the POWER method, specific hyperparameters are fixed as follows: trade-off parameter $\alpha$ at 0.5, number of experience nodes per task $b$ at 1, and the K-Nearest Neighbor parameter at 1, as per Eqs. (3), (5), and (LABEL:eq:_knn), respectively. Exploratory hyperparameters include the the threshold distance $\eta$ set within $\{0.01,0.1,0.5\}$ in Eq. (4), trade-off parameter $\beta$ in Eq. (8) and the past task decay coefficient $\phi$ varied from 0 to 1 in increments of $0.01$ in Eq. (13). Finally, the pseudo prototype reconstruction process in Eq. (12) is optimized with 300 iterations using the LBFGS optimizer [65], which has an initial learning rate of 1.0. Results are reported as the mean and variance across 10 standardized runs.
Experiment Environment. The experimental machine is an Intel(R) Xeon(R) Gold 6240 CPU @ 2.60GHz and NVIDIA A100 with 80GB memory and CUDA 12.4. The operating system is Ubuntu 22.04.5 with 251GB memory.
TABLE II: Performance comparison of POWER and baselines, where the best and second results are highlighted in bold and underline.
Dataset Cora CiteSeer OGB-arxiv Computers Physics Squirrel Flickr Roman-empire
Methods AM $\uparrow$ FM $\downarrow$ AM $\uparrow$ FM $\downarrow$ AM $\uparrow$ FM $\downarrow$ AM $\uparrow$ FM $\downarrow$ AM $\uparrow$ FM $\downarrow$ AM $\uparrow$ FM $\downarrow$ AM $\uparrow$ FM $\downarrow$ AM $\uparrow$ FM $\downarrow$
FedAvg [33] 40.14 ±1.01 78.92 ±1.84 42.50 ±0.73 59.30 ±1.11 21.90 ±0.99 57.57 ±1.54 19.81 ±0.07 94.94 ±1.71 60.63 ±1.80 75.98 ±3.58 32.33 ±0.95 59.03 ±2.67 27.46 ±2.00 37.35 ±3.02 17.19 ±0.92 42.11 ±1.60
FedSage+ [26] 30.41 ±2.39 87.63 ±0.28 33.70 ±4.51 58.76 ±6.29 15.22 ±0.83 56.87 ±1.21 12.22 ±1.77 95.19 ±2.35 58.24 ±2.18 79.82 ±3.04 29.52 ±2.18 79.82 ±3.04 25.58 ±2.54 36.12 ±2.07 13.35 ±1.33 40.27 ±3.15
Fed-PUB [40] 30.02 ±1.28 90.88 ±0.35 33.48 ±2.34 73.66 ±2.38 15.17 ±0.15 56.94 ±0.74 11.93 ±1.54 95.46 ±0.46 58.75 ±0.62 79.72 ±1.24 28.82 ±0.42 43.52 ±0.56 25.47 ±1.10 36.27 ±1.27 11.56 ±0.20 41.60 ±1.02
FedGTA [41] 38.62 ±0.66 80.99 ±1.31 47.00 ±0.84 54.05 ±1.65 16.80 ±0.62 62.89 ±1.83 19.48 ±0.60 96.40 ±3.35 61.62 ±2.07 74.10 ±4.09 30.84 ±0.38 42.47 ±2.29 27.78 ±1.82 40.05 ±1.74 14.16 ±1.29 46.54 ±0.66
Fed-TWP [21] 40.34 ±1.10 78.67 ±1.63 42.38 ±1.33 58.94 ±1.48 22.44 ±0.35 57.31 ±1.42 19.02 ±1.81 94.11 ±3.65 61.43 ±2.22 74.38 ±4.45 32.34 ±0.55 57.08 ±3.84 30.43 ±4.20 37.41 ±6.37 17.58 ±0.86 42.02 ±1.36
Fed-ERGNN [20] 60.40 ±2.78 45.68 ±4.20 49.92 ±2.02 39.19 ±2.41 32.63 ±3.48 48.78 ±5.75 39.79 ±7.00 20.27 ±13.82 85.23 ±2.74 26.02 ±5.56 35.13 ±2.12 41.85 ±4.38 31.79 ±3.93 22.52 ±5.85 24.71 ±0.87 27.87 ±3.56
Fed-DSLR [23] 61.21 ±3.25 43.77 ±5.10 49.55 ±2.16 40.92 ±3.27 31.69 ±4.32 50.32 ±3.22 40.32 ±5.30 18.56 ±8.31 84.75 ±5.74 28.12 ±6.56 35.21 ±3.15 41.33 ±5.27 31.55 ±2.47 24.18 ±6.30 24.32 ±2.87 28.22 ±4.13
GLFC [27] 52.04 ±5.34 55.80 ±7.13 46.30 ±3.27 51.24 ±4.35 30.27 ±4.34 52.43 ±5.25 38.43 ±7.62 23.56 ±3.27 67.01 ±4.87 60.97 ±4.22 35.99 ±6.78 40.21 ±4.28 30.80 ±3.74 35.27 ±6.03 16.86 ±3.07 41.52 ±4.85
TARGET [29] 51.64 ±4.18 56.69 ±2.37 44.62 ±2.14 56.95 ±3.15 30.52 ±5.19 53.02 ±4.20 37.75 ±1.54 26.82 ±1.79 75.48 ±2.06 19.88 ±3.18 32.75 ±2.03 51.59 ±3.11 31.07 ±3.24 25.12 ±2.20 14.37 ±1.11 45.02 ±1.95
LANDER [30] 52.85 ±5.07 53.57 ±3.24 45.93 ±3.20 52.45 ±4.33 30.45 ±4.12 54.31 ±3.73 38.13 ±2.64 24.27 ±3.28 73.24 ±4.51 22.46 ±4.20 33.21 ±3.35 52.03 ±5.17 30.58 ±4.12 28.32 ±3.27 15.42 ±2.35 44.27 ±3.53
POWER (Ours) 65.74 ±5.11 28.10 ±4.95 54.47 ±5.21 28.94 ±3.19 36.19 ±3.17 30.52 ±4.22 43.26 ±2.94 12.59 ±3.03 89.17 ±2.54 13.03 ±4.53 40.27 ±3.22 35.24 ±2.98 35.79 ±4.15 18.13 ±3.24 26.54 ±2.35 16.23 ±2.26
V-B Performance Comparison (Answer for Q1)
To answer Q1, we present a performance comparison of our proposed POWER framework against various baselines in Table. II. As shown, POWER consistently outperforms all baselines across both AM and FM metrics. We provide a detailed discussion below for each baseline category, explaining the reasons for their sub-optimal performance.
Comparison to FL/FGL Fine-tuning. These methods lack a targeted design for effectively alleviating LGF across sequential tasks, leading to notably poor retention of prior task performance. Consequently, they consistently exhibit both low AM and high FM. Particularly, algorithms initially designed for FGL (i.e., FedSage+, Fed-PUB, and FedGTA) frequently perform poorly compared to the simpler FedAvg. This phenomenon stems from their inherently complex designs, which focus on learning local data and even tend toward overfitting them, thereby further exacerbating the forgetting of previous tasks when each client’s local graph evolves.
Comparison to Federated CGL. These methods simply integrate the centralized CGL algorithm with FedAvg, enabling them to mitigate LGF but not effectively address GEC. As observed, despite these methods retaining a certain degree of memory for previous tasks, they still exhibit a notable performance gap compared to POWER (e.g., comparison between Fed-ERGNN and POWER on the Cora dataset), which is primarily caused by GEC, as outlined in our insights in Sec. III. Additionally, a counterintuitive but important observation emerges: the performance of FedTWP remains on par with FedAvg in most cases, almost entirely losing the ability to retain the knowledge of older tasks. We attribute this issue to its memory mechanism for previous tasks, which primarily relies on adding regularization terms to a subset of parameters sensitive to changes in prior task loss during training new tasks. However, under the FCGL setting, the local model parameters are consistently replaced by the global model in each round, severely disrupting the parameter-preserving process of FedTWP, thus leading to significant LGF.
Figure 4: Convergence curves for the proposed POWER framework and six baseline methods across eight benchmark graph datasets observed during training on the final task, where “Round 0” corresponds to the state prior to initiating training on the final task.
Comparison to Federated Continual Learning for CV. These methods are specifically tailored for federated continuous learning in computer vision contexts. However, they exhibit noticeable shortcomings in the FCGL scenario. Although they exhibit clear superiority over most fine-tuning methods, their performance is still significantly lacking when compared with Fed-ERGNN and POWER. This stems from their excessive reliance on image data augmentation strategies or the necessity for meticulous training of generative models in the visual domain to produce pseudo images. In the FCGL setting, merely removing the image data augmentation strategies or employing a simple GNN-based generative model does not yield satisfactory results. Additionally, these methods also fail to incorporate designs for effectively addressing GEC.
V-C Efficient Analysis (Answer for Q2)
To address Q2, we provide a detailed analysis of convergence speed and communication overhead, as outlined below.
Convergence Speed. We present the convergence curves of POWER alongside six continual learning-related baselines: Fed-TWP, Fed-ERGNN, Fed-DSLR, GLFC, TARGET, and LANDER. Notably, methods such as FL/FGL Fine-tuning are excluded from this comparison, as their limitations in FCGL scenarios have already been established in Q1. The results, illustrated in Fig. 4, demonstrate that POWER consistently converges faster than all baselines, showcasing its enhanced efficiency. This efficiency is critical for real-world applications, as it requires fewer communication rounds to achieve optimal performance, ultimately reducing communication overhead significantly. We attribute this to the effectiveness of the pseudo prototype reconstruction module in capturing localized expertise from each client. Additionally, the trajectory-aware knowledge transfer mechanism enables the global model to assimilate this expertise more effectively. These factors contribute to POWER’s ability to achieve fast convergence.
TABLE III: Theoretical communication overhead comparison of POWER and FedAvg. Notably, FedAvg represents the lowest communication cost of mainstream federated learning processes.
Method Content Overhead
FedAvg GNN Parameters $\mathcal{O}(KTRN_{\mathbf{\Theta}_{\text{GNN}}})$
POWER (Ours) GNN Parameters Prototype Gradients $\mathcal{O}(KTRN_{\mathbf{\Theta}_{\text{GNN}}}+$ $KCN_{\mathbf{\Theta}_{\text{GE}}})$
Communication Overhead. To assess POWER’s practicality in real-world settings, we compare its theoretical communication cost with the baseline FedAvg, which also underpins Fed-ERGNN, Fed-TWP, and Fed-DSLR, representing the minimal communication cost among mainstream federated learning approaches. Table. III summarizes the results, where $K,T,R,C$ represent the number of clients, tasks, communication rounds, and total classes across all tasks, respectively. Additionally, $N_{\mathbf{\Theta}_{\text{GNN}}}$ and $N_{\mathbf{\Theta}_{\text{GE}}}$ denote the parameter counts for the local GNN and gradient encoding network. Both POWER and FedAvg require transmitting GNN parameters to the server, incurring a communication cost of $\mathcal{O}(KTRN_{\mathbf{\Theta}_{\text{GNN}}})$. POWER introduces an additional term for the prototype gradient, sent from clients to the server, with a cost of $\mathcal{O}(KCN_{\mathbf{\Theta}_{\text{GE}}})$. Importantly, in practical settings, $\mathcal{O}(KCN_{\mathbf{\Theta}_{\text{GE}}})$ is typically far smaller than $\mathcal{O}(KTRN_{\mathbf{\Theta}_{\text{GNN}}})$. This reduction can be attributed to several natural factors: (1) Real-world graph data tends to be complex and noisy, often requiring a large number of communication rounds $R$ for convergence, while POWER only uploads prototype gradients in the first round of communication for each task. (2) Due to the personalized nature of local data sources and collection methods, each client typically lacks data for some categories and, as a result, does not transmit prototypes for all classes $C$. (3) The GNN models in practice have significantly larger parameter counts compared to our lightweight gradient encoding network (i.e., $N_{\mathbf{\Theta}_{\text{GE}}}\ll N_{\mathbf{\Theta}_{\text{GNN}}}$).
Figure 5: Sensitivity analysis for the replay buffer size $b$ across eight benchmark graph datasets.
TABLE IV: Ablation study results on the Cora dataset, with the complete POWER framework as the baseline highlighted in blue.
POWER Components Cora
LGF Module GEC Module AM $\uparrow$ FM $\downarrow$
[…]
V-D Ablation Study (Answer for Q3)
To address Q3, we explore the composition of POWER, with a focus on two key modules: (1) LGF module, which integrates local-global coverage maximization, i.e., Eqs. (3), (4), and (5) alongside experience node replay, i.e., Eq. (8); and (2) GEC module, incorporating pseudo prototype reconstruction, i.e., Eqs. (10), (11), (12), and (13), and trajectory-aware knowledge transfer, i.e., Eqs. (LABEL:eq:_knn) and (IV-B). To systematically evaluate the contribution of each component, we conduct a comprehensive ablation study using the Cora dataset, analyzing the performance of POWER under configurations where only one module is activated. Furthermore, we delve into variations within each module. Specifically, for the LGF module, we adopt local coverage maximization instead of the combined local-global approach (Local CM for short); for the GEC module, we replace the cumulative label distribution with a non-cumulative alternative (Non-cumulative for short). The results of the ablation study are presented in Table. IV.
Inter-module Ablation. The removal of the LGF Module leads to significant performance declines of approximately 22% and 46% in the AM and FM metrics, respectively, underscoring the critical role of the LGF challenge in influencing FCGL’s effectiveness. Conversely, the exclusion of the GEC Module results in reductions of 4% and 14%. Although not negligible, it is less of a loss than removing LGF, suggesting that the GEC challenge only limits further performance enhancements. Finally, the simultaneous removal of both modules reduces POWER to a mere FedAvg algorithm.
Intra-module Ablation. Both the Local CM and Non-cumulative configurations exhibit performance declines, affirming that our local-global coverage maximization strategy benefits from incorporating global insights to select more representative experience nodes for past tasks. Additionally, by leveraging the cumulative label distribution, POWER can effectively capture the graph evolution trajectory, taking into account the past task knowledge acquired by each local GNN.
In summary, both these modules and their designed components play a pivotal role in shaping the overall performance.
V-E Sensitivity Analysis (Answer for Q4)
Figure 6: Sensitivity analysis for various combinations of the trade-off parameters $\beta$ and the decay coefficient $\phi$ on the Cora dataset.
To address Q4, we conduct a sensitivity analysis of the hyperparameters in the proposed POWER framework.
First, we investigate the impact of the replay buffer size $b$ as defined in Eq. (5), with findings depicted in Fig. 5. Increasing $b$ improves both AM and FM. Notably, our default configuration is set to $b=1$ to minimize storage costs, as larger $b$ values increase the local storage required to retain past nodes. However, in scenarios with ample storage capacity, $b$ can be appropriately improved to obtain better prediction performance.
Next, we analyze the sensitivity of POWER to the trade-off parameters $\beta$ in Eq. (8) and the decay coefficient $\phi$ in Eq. (13), both of which affect how past task knowledge is retained. Fig. 6 presents the experimental results on the Cora dataset, focusing on values near the identified optimal range. As observed, a large $\beta$ may lead to insufficient replaying for old task knowledge, while a small $\phi$ might fail to account for it adequately when assessing each local GNN’s expertise. Notably, the proposed POWER framework demonstrates stable performance in the vicinity of suitable value combinations for $\beta$ and $\phi$, maintaining holistic robustness.
Figure 7: Performance under sparse settings for the proposed POWER framework and six baseline methods on the Physics dataset.
V-F Performance under Sparse Settings (Answer for Q5)
In response to Q5, we conducted experiments to assess the resilience of the POWER framework relative to six baselines across various sparsity conditions: (1) feature sparsity, where node features are selectively obscured to mimic incomplete information; (2) edge sparsity, which simulates partial graph connectivity by randomly removing edges; (3) label sparsity, where only a subset of nodes have available labels to represent scenarios with incomplete supervision; and (4) client sparsity, characterized by limited client participation in federated training rounds due to factors such as network constraints or hardware limitations. These four sparsity settings reflect the natural limitations encountered in real-world FCGL scenarios.
Sparse Feature/Edge/Label. The results for sparse feature/edge/label settings are depicted in Fig. 7, where panels (a)-(c) present the AM metric, and (e)-(g) present the FM metric. As observed, POWER reveals superior performance across all degrees of sparsity. We attribute this superiority to its pseudo prototype reconstruction mechanism and efficient knowledge transfer processes. Specifically, for feature sparsity, POWER calculates prototypes using node features within the same category, effectively compensating for missing features by utilizing the complementary data from multiple nodes. For edge sparsity, although POWER does not specifically target missing connections, its superior performance likely stems from the strong overall learning architecture and effective knowledge transfer process. For label sparsity, POWER’s server-side knowledge transfer process allows each client to share its local insights in a privacy-preserving manner, which enhances the global model’s ability to compensate for sparse local supervision, effectively learning from limited label information.
Sparse Client Participation. The experimental results for client sparsity settings are depicted in Fig. 7, where panel (d) presents the AM metric, and (h) presents the FM metric. As observed, POWER maintains stable performance across various participation ratios. We attribute this robustness in client-sparse settings to its effective prototype reconstruction mechanism, detailed in the methods section. Specifically, when POWER receives prototype gradients from a participating client, it effectively reconstructs and stores these prototypes in a global buffer. Thus, even if a client discontinues participation in subsequent federation rounds, the knowledge from its prototypes continuously contributes to the knowledge transfer process, thereby optimizing the global model.
In summary, in the FCGL context, POWER outperforms traditional CGL methods and vision-focused federated learning algorithms, demonstrating superiority for sparsity challenges.
VI Conclusion
This paper pioneers the exploration of federated continual graph learning (FCGL) for node classification, bridging the gap between the idealized experimental setups in previous centralized continual graph learning (CGL) studies and the challenges encountered in real-world decentralized deployments. We introduce a practical solution to advance graph continual learning by harnessing collective intelligence across distributed graph databases. Through empirical analysis, we investigate the data characteristics, feasibility, and effectiveness of FCGL, identifying two critical challenges: local graph forgetting (LGF) and global expertise conflict (GEC). To address these challenges, we propose a novel FCGL framework that selectively replays experience nodes with maximum local-global coverage to mitigate the LGF, while employing a pseudo-prototype reconstruction process and a trajectory-aware knowledge transfer technique to resolve GEC. A promising avenue for future research involves developing more effective methods to identify compatibility and conflict relationships of expertise across evolving client subgraphs, facilitating improved global aggregation. This could further minimize the forgetting of old tasks and enhance adaptation to new tasks, advancing the potential of FCGL in dynamic environments.